OpenStudy (anonymous):

f(x)=6-5sin2(x-pi/3)

OpenStudy (anonymous):

state domain and range

OpenStudy (anonymous):

Domain: All real numbers Range: -6<f(x)<6, this could also be all real numbers.

OpenStudy (anonymous):

how?

OpenStudy (dls):

-1<5sin2(x-pi/3)<1 5<6-5sin2(x-pi/3)<7

OpenStudy (dls):

5 and 7 are in open intervals

OpenStudy (anonymous):

the solution sheet says D:(-oo,oo) R: [1,11]

OpenStudy (dls):

oops sorry

OpenStudy (dls):

6-5sin2(x-pi/3)..................... -1<=sin(x-pi/3)<=1 -5<=5sin(x-pi/3)<=5 ( on multiplying by 5) 5>=-5sin(x-pi/3)>=-5 ( on multiplying with -1) 11>=6-5sin(x-pi/3)>=1 (on adding 6)